OpenStudy (anonymous):

Don't know how to start???? (proving identity) 1+TAN theta Tan 2theta = TAN 2theta cot theta-1 Please HELP?

OpenStudy (midhun.madhu1987):

can you draw the question or post the screenshot of question?

OpenStudy (anonymous):

i'll try

OpenStudy (anonymous):

\[1+ \tan \theta \tan 2\theta = \tan 2\theta \cot \theta - 1\]

OpenStudy (anonymous):

im not sure if i start on the left side or the right side? as far as trying to transform it?

OpenStudy (midhun.madhu1987):